# Visual Evolutionary Optimization on Combinatorial Problems with Multimodal Large Language Models: A Case Study of Influence Maximization

### v1

@@ Abstract
Graph-structured combinatorial problems in complex networks are prevalent in many domains, and are computationally demanding due to their complexity and non-linear nature. Traditional evolutionary algorithms (EAs), while robust, often face obstacles due to content-shallow encoding limitations and lack of structural awareness, necessitating hand-crafted modifications for effective application. In this work, we introduce an original framework, Visual Evolutionary Optimization (VEO), leveraging multimodal large language models (MLLMs) as the backbone evolutionary optimizer in this context. Specifically, we propose a context-aware encoding way, representing the solution of the network as an image. In this manner, we can utilize MLLMs’ image processing capabilities to intuitively comprehend network configurations, thus enabling machines to solve these problems in a human-like way. We have developed MLLM-based operators tailored for various evolutionary optimization stages, including initialization, crossover, and mutation. Furthermore, we propose that graph sparsification can effectively enhance the applicability and scalability of VEO on large-scale networks, owing to the scale-free nature of real-world networks. We demonstrate the effectiveness of our method using a well-known task in complex networks, influence maximization, and validate it on eight different real-world networks of various structures. The results have confirmed VEO’s reliability and enhanced effectiveness compared to traditional evolutionary optimization.
@@ I Introduction
Note that the choice of encoding scheme plays a crucial role in solving discrete problems, as it determines how solutions are represented and manipulated within the evolutionary optimization framework. Different discrete tasks require tailored encoding strategies in EAs, such as permutation and path encodings for scheduling [9] and routing problems [10] or binary and label-based (node or edge index) encodings for subset selection problems, such as influence maximization [11] and community deception [12]. However, these traditional encodings represent nodes or paths in an abstract manner, failing to capture their roles within the network. As a result, evolutionary operators like crossover and mutation are applied without an inherent understanding of the underlying entities or relationships. This lack of structural awareness can lead to naive modifications that may not respect the structural properties of the graph.
Multimodal large language models (MLLMs) are extensively trained on diverse datasets, enabling them to effectively process multiple data types with inherent adaptability. [13, 14]. This capability presents a unique opportunity to enhance evolutionary optimization fundamentally. In this work, we introduce a novel encoding methodology that visually represents solutions as images. By leveraging this visual encoding strategy, we unlock the potential of MLLMs in evolutionary optimization and propose an innovative framework called Visual Evolutionary Optimization (VEO). This framework enables the model to analyze and manipulate solutions in a way that mirrors human decision-making in evolutionary operations, such as crossover and mutation. The motivations and advantages behind VEO operate on two levels and are discussed as follows:
@@
  • Scalability and Inference Efficiency: Image representations maintain a consistent input size regardless of graph complexity, avoiding issues like token explosion in large-scale networks. This characteristic ensures efficient inference and improved scalability within MLLM frameworks.
(2) Optimization-level: MLLMs as Evolutionary Operators
  • Human-Like Visual Reasoning: MLLMs can interpret the structure of graph-encoded images and reason about solution quality in a human-like manner, enabling high-quality decision-making for offspring generation within the EA framework without relying on random search or complex heuristics.
  • Generalization and Adaptability: MLLMs adapt across diverse and dynamic graph structures without requiring problem-specific algorithm design. They flexibly respond to changes in topology or constraints, offering broad applicability and real-time adaptability.
@@
Effective visualization is essential for utilizing VEO on graph-structured problems, particularly when dealing with large networks. Plotting all nodes on a limited canvas often results in cluttered visuals, obscuring critical structural details and connections. Node labels, essential for MLLM-based operations like initialization, crossover, and mutation, must remain clear and readable to ensure that MLLMs can accurately identify each node’s role and relationship within the network. As noted in [15], many real-world networks exhibit a scale-free structure, where a small subset of nodes holds disproportionately high centrality or importance. By concentrating on these key nodes and reducing overall graph complexity without significantly compromising solution quality, we are able to address large-scale, real-world problem instances more effectively. To this end, we propose a community-level graph sparsification method that compresses large networks into more manageable forms while preserving essential structural information necessary for meaningful analysis and optimization.
@@
$\bullet$ Our paper introduces a new context-aware encoding scheme for graph-structured combinatorial problems, depicting solutions as images that capture extensive structural details of networks. The proposed representation allows MLLMs to execute evolutionary operations with a profound understanding of the network’s topology. We then develop MLLM-based operators tailored for various stages of evolutionary optimization, including initialization, crossover, and mutation, revolutionizing how solutions are manipulated within an evolutionary framework.
$\bullet$ We explore how to effectively improve the applicability and scalability of MLLMs by leveraging graph sparsification. Specifically, we propose a community-level sparsification method that reduces the size of large-scale networks while preserving key structural information. As part of this process, a new community merging technique is employed to consolidate close communities, enhancing sparsification efficiency while maintaining meaningful graph semantics.
$\bullet$ We have comprehensively investigated the reliability of VEO, such as the validity and effectiveness of MLLMs output. Several key parameters such as the influence of the number of nodes in the plotted network as well as the different-level tasks on the quality of MLLMs are also studied. With the illustration of a well-known task, i.e., influence maximization, the experiments are conducted on eight different real-world networks and the ablation results suggest the proposed VEO can greatly facilitate different stages of evolutionary optimization.
@@
The structure of this paper is as follows: Section II discusses previous research on evolutionary optimization, LLMs and combinatorial problems. Section III details the proposed MLLM-driven evolutionary optimization. In Section IV, we evaluate the effectiveness of VEO and analyze various parameters. The future work for the MLLM-based evolutionary optimization on graph-structured combinatorial problems is discussed in Section V. Finally, the paper concludes with Section VI.
@@ II-B LLMs-Assisted Evolutionary Optimization
Recently, Wu et al. [36] classified existing works regarding the synergy of LLMs and evolutionary computation into two main branches: LLMs-based search operators [37, 38] and LLMs-based optimization algorithm automation [39, 40, 41]. Our study can be roughly classified into the first category. In [42], LLMs are employed as crossover operators to derive new solutions from parental inputs. Brownlee [43] also presented LLMs effectively functioning as mutation operators that enhance the search process. Liu et al. [44] introduced a novel framework known as LLMs-driven EA (LMEA), which utilizes LLMs for both crossover and mutation operations. Furthermore, LLMs-based search operators can also be adapted for multi-objective scenarios by segmenting traditional optimization tasks into sub-problems [45]. The related application can be found in [46], where protein properties are optimized with LLMs-enhanced evolutionary operators.
@@ II-C LLMs on Graph-Structured Problems
The recent advancements in LLMs have led to significant progress across diverse fields including sentiment analysis [47], optimization [48], and social sciences [49]. This raises the question of whether the capabilities of LLMs can extend to graph-related tasks [50, 51]. However, as indicated by studies such as [52, 53], LLMs struggle with graph-structured data and often fail to achieve satisfactory results on fundamental tasks, especially as graph complexity increases.
To date, there are also some explorations of the applications of MLLMs in graph tasks. For example, Wei et al. [54] converted graphs to images and feeds them into MLLMs for seven fundamental graph reasoning tasks, showcasing the advantages of using MLLMs with image representations of graphs with its visual intelligence. Similarly, VisionGraph [55] explored leveraging large multimodal models for graph theory problems in a visual context, establishing a toolchain for eight complex graph problem tasks. Beyond the basic graph problems, research on MLLMs for combinatorial challenges is emerging. Huang et al. [56] have combined visual and textual data to address the traveling salesman problem (TSP). Subsequently, Elhenawy et al. [57] focused solely on graphical data to determine optimal routing, assessing various MLLMs.
As far as we know, the synergy of MLLMs and evolutionary optimization is still underexplored. In this work, we will explore how MLLMs facilitate evolutionary optimization in complex network scenarios with a case study of influence maximization.
@@ III-B Graph Sparsification
Consider a graph $G=(V,E)$ with a set of communities $\widetilde{\mathcal{C}}=\{\mathcal{C}_{1},\mathcal{C}_{2},\ldots,\mathcal{C}_{k}\}$, where each $\mathcal{C}_{i}\subset V$ denotes a cluster of nodes. The simplified network should retain key nodes and connections, preserving the core community structure of the original network. The proposed sparsification process is as follows (see Algorithm 1 for the pseudocode): We begin by determining selected nodes from each community, that is
@@
$V^{\prime}=\bigcup_{i=1}^{m}\{v_{j}\in C_{i}\mid b(v_{j})\geq b(v_{l}),\forall v_{l}\in C_{i}\setminus V^{\prime}\},$ (5)
where $b(v)$ refers to the betweenness value of node $v$ and $|V^{\prime}\cap C_{i}|=|\mathcal{C}^{\prime}_{i}|$.
@@
In addition, any isolated nodes will also be removed, and only the largest connected component is retained. The node and edge sets of refined network are updated as:
$V^{\prime}$ =$=\{v\in V^{\prime}\mid d(v)>0\},$ (7)
$E^{\prime}$ =$=\{(u,v)\in E^{\prime}\mid u,v\in V^{\prime}\}.$
@@
2:Identify the community structure $\mathcal{C}=\{\mathcal{C}_{1},\mathcal{C}_{2},\ldots,\mathcal{C}_{m}\}$
3:for each community $\mathcal{C}_{i}$ in $\mathcal{G}$ do
@@
8:Construct $V^{\prime}\leftarrow\bigcup_{i=1}^{k}V^{\prime}_{i}$
9:$E^{\prime}\leftarrow\{(u,v)\in E:u\in V^{\prime}\text{ and }v\in V^{\prime}\}$
@@ III-C Community Merging
Most existing community detection techniques often lead to clusters with highly imbalanced sizes, including many small or even single-node communities that contribute little to the overall network structure. This imbalance can undermine the effectiveness of the aforementioned graph sparsification, indicating the importance of merging these small-scale communities. Given a graph $G=(V,E)$ with an initial community structure $\mathcal{C}$, where each community $C_{i}\subseteq V$, our goal is to reduce the total number of communities to a target number $N_{\mathcal{C}}^{*}$. We define the size of each community $C_{i}$ as $|C_{i}|$, and the algorithm proceeds while the number of updated communities $|\widetilde{\mathcal{C^{\prime}}}|$ exceeds $N_{\mathcal{C}}^{*}$. The community merging process involves several key steps:
@@
    Size Identification: Identify the smallest community $C_{\text{min}}$, where $|C_{\text{min}}|=\min\{|C_{i}|\}$ for all $i$.
@@
    Connectivity Analysis: For each edge $(v_{i},v_{j})\in E$, increment an edge count between communities for edges where $v_{i}\in C_{\text{min}}$ and $v_{j}\in C_{k}$, and vice versa. This edge count $e(C_{\text{min}},C_{k})$ measures the connectivity between $C_{\text{min}}$ and each other community.
@@
    Determination of Closest Community: Identify the community $C_{\text{close}}$ with the maximum edge count to $C_{\text{min}}$, i.e.,
    $C_{\text{close}}=\arg\max_{C_{k}}e(C_{\text{min}},C_{k}).$
After each merging operation, the community structure $\mathcal{C}$ is updated, and the process repeats until the number of communities equals $N_{\mathcal{C}}^{*}$. The process of community merging can be found in Algorithm 2.
@@
5:     for each community $\mathcal{C}_{k}\in\mathcal{C}$ such that $\mathcal{C}_{k}\neq\mathcal{C}_{\min}$ do
@@
13:     Merge $\mathcal{C}_{\min}$ and $\mathcal{C}_{\text{close}}$ to obtain $\mathcal{C}_{\text{merged}}$
@@ III-D MLLM-based Evolutionary Optimizer
When solving graph-structured combinatorial problems, traditional EAs typically utilize list-based encodings, consisting of sequences of node indices. While this approach is straightforward, it has some inherent drawbacks. Crossover and mutation operations could be compromised due to the absence of spatial awareness concerning the network’s structure.
@@
5:$\mathcal{P}\leftarrow$ Initialize population with MLLMs by encoding $G$ into images
@@
[…]
10:     Apply MLLM-based mutation to images of Offspring to alter node sets
@@
13:$\mathcal{S^{\prime}}\leftarrow$ Select the solution with the highest fitness in $\mathcal{P}$
@@
Figure 1: The diagram of MLLM-driven VEO and the comparison with traditional evolutionary optimization, illustrated with influence maximization. (a): The original network presents a graph with nodes and edges, where two solutions (seed sets) are shown: Solution 1 [A, C, P] and Solution 2 [A, E, H], representing selected nodes to maximize influence spread. (b): Traditional encoding and reproduction represent solutions as index lists. However, the crossover process often generates invalid solutions, such as duplicate nodes (e.g., [A, A, P]) or suboptimal combinations (e.g., [A, E, P] where nodes A and E are low-degree). (c): Our scheme encodes the solution in the form of an image where seed nodes of the solution are distinctly highlighted. Crossover intelligently combines nodes from parent solutions to produce valid offspring filter the low-degree nodes (e.g., nodes A and P), while mutation replaces ineffective nodes with more suitable ones to enhance influence spread (replace node E with F). This approach ensures validity and better solution quality. (d): Prompt for reproduction provides instructions for an MLLM: in crossover, the MLLM examines network images and suggests an optimal combination of seed nodes; in mutation, it identifies ineffective nodes for replacement by removing nodes with minimal influence and proposing better non-seed nodes (new).
TABLE I: Structure of prompts for MLLM-based evolutionary operators of different phases. The prompt begins with context-setting, which introduces the input information and clarifies the role of the agents. The latter part of the prompt specifies the desired output format and includes any necessary guidance or restrictions.
@@
To overcome these challenges, we propose a new context-aware framework VEO that incorporates visual encoding and direct interaction with network images by utilizing the capabilities of MLLMs. This method transforms the network’s solution into an image format, enhancing machines’ understanding of how specific nodes influence particular tasks. The diagram of our method is shown in Figure 1 and the process of MLLM-based evolutionary optimization is in Algorithm 3.
@@
which produces an image of the effective graph $G$ using the layout $\Theta$ via visualization software. In contrast, to encode specific candidate solutions, we define the function
@@
MLLM-Based Mutation: A mutation operator $M(\cdot)$ is similarly applied on the network, yielding:

### v2

# Visual Evolutionary Optimization on Graph-Structured Combinatorial Problems with MLLMs: A Case Study of Influence Maximization

@@ Abstract
Graph-structured combinatorial problems in complex networks are prevalent in many domains, and are computationally demanding due to their complexity and non-linear nature. Traditional evolutionary algorithms (EAs), while robust, often face obstacles due to content-shallow encoding limitations and lack of structural awareness, necessitating hand-crafted modifications for effective application. In this work, we introduce an original framework, visual evolutionary ptimization (VEO), leveraging multimodal large language models (MLLMs) as the backbone evolutionary optimizer in this context. Specifically, we propose a context-aware encoding way, representing the solution of the network as an image. In this manner, we can utilize MLLMs’ image processing capabilities to intuitively comprehend network configurations, thus enabling machines to solve these problems in a human-like way. We have developed MLLM-based operators tailored for various evolutionary optimization stages, including initialization, crossover, and mutation. Furthermore, we propose that graph sparsification can effectively enhance the applicability and scalability of VEO on large-scale networks, owing to the scale-free nature of real-world networks. We demonstrate the effectiveness of our method using a well-known task in complex networks, influence maximization, and validate it on eight different real-world networks of various structures. The results have confirmed VEO’s reliability and enhanced effectiveness compared to traditional evolutionary optimization.
@@ I Introduction
The choice of encoding scheme plays a crucial role in solving discrete problems, as it determines how solutions are represented and manipulated within the evolutionary optimization framework. Different discrete tasks require tailored encoding strategies in EAs, such as permutation and path encodings for scheduling [9] and routing problems [10], or binary and label-based (node or edge index) encodings for subset selection problems, such as influence maximization [11] and community deception [12]. However, these traditional encodings represent nodes or paths in an abstract manner, failing to capture their roles within the network. As a result, evolutionary operators like crossover and mutation are applied without an inherent understanding of the underlying entities or relationships. This lack of structural awareness can lead to naive modifications that may not respect the structural properties of the graph.
Multimodal large language models (MLLMs) are extensively trained on diverse datasets, enabling them to effectively process multiple data types with inherent adaptability [13, 14]. This capability presents a unique opportunity to enhance evolutionary optimization fundamentally. In this work, we introduce a novel encoding methodology that visually represents solutions as images. By leveraging this visual encoding strategy, we unlock the potential of MLLMs in evolutionary optimization and propose an original framework called Visual Evolutionary Optimization (VEO). This framework enables the model to analyze and manipulate solutions in a way that mirrors human decision-making in evolutionary operations, such as crossover and mutation. The motivations and advantages behind VEO operate on two levels as follows:
@@
  • Scalability and Inference Efficiency: Image-based representations maintain a consistent input size regardless of graph complexity, thereby avoiding issues such as token explosion, which commonly arises when representing large-scale networks in textual form. This characteristic enables more efficient inference and enhances scalability within MLLM frameworks.
(2) Optimization-level: MLLMs as Evolutionary Optimizers
  • Human-Like Visual Reasoning: MLLMs can interpret the structure of image-encoded graph and reason about solution quality in a human-like manner, enabling high-quality decision-making for offspring generation within the EA framework without relying on random search or complex heuristics.
  • Generalization and Adaptability: MLLMs can adapt across diverse and dynamic graph structures without requiring problem-specific algorithm design. They can also flexibly respond to changes in topology, offering broad applicability and real-time adaptability.
@@
Effective visualization is essential for utilizing VEO on graph-structured problems, particularly when dealing with large networks. Plotting all nodes on a limited canvas often results in cluttered visuals, obscuring critical structural details and connections. Node labels, essential for MLLM-based operations like initialization, crossover, and mutation, must remain clear and readable to ensure that MLLMs can accurately identify each node’s role and relationship within the network. As noted in [15], many real-world networks exhibit a scale-free structure, where a small subset of nodes holds disproportionately high centrality or importance. By concentrating on these key nodes and reducing overall graph complexity without significantly compromising solution quality, we are able to address large-scale problem instances more effectively. To this end, we propose a community-level graph sparsification method that compresses large networks into more manageable scales while preserving essential structural information necessary for meaningful analysis and optimization.
@@
$\bullet$ Our paper presents a novel context-aware encoding scheme for graph-structured combinatorial optimization, where solutions are represented as images that encapsulate rich structural information of the underlying networks. The proposed representation allows us to further develop evolutionary operations with a profound understanding of the network’s topology. More precisely, we introduce MLLM-based operators tailored to various stages of evolutionary optimization, including initialization, crossover, and mutation, fundamentally redefining how solutions are generated, manipulated, and refined within the evolutionary framework.
$\bullet$ We explore how to effectively improve the applicability and scalability of MLLMs by leveraging graph reduction techniques. Specifically, we propose a community-level sparsification method that reduces the size of large-scale networks while preserving key structural information. As part of this process, a new community merging technique is employed to consolidate close communities, enhancing sparsification efficiency while maintaining meaningful graph semantics.
$\bullet$ We have comprehensively investigated the reliability of VEO, such as the validity and effectiveness of MLLMs output. Several key parameters such as the influence of the number of nodes in the plotted network as well as the different-level tasks on the quality of MLLMs are also studied. The experiments are conducted on eight different real-world networks and the ablation results suggest the proposed VEO can greatly facilitate different stages of evolutionary optimization.
@@
The structure of this paper is as follows: Section II discusses previous research on evolutionary optimization, LLMs and combinatorial problems. Section III details the proposed MLLM-driven evolutionary optimization. In Section IV, we evaluate the effectiveness of VEO and analyze various parameters. The future work for the MLLM-based evolutionary optimization on graph-structured combinatorial problems is discussed in Section V. Finally, Section VI presents the conclusions of the paper.
@@ II-B LLMs-Assisted Evolutionary Optimization
Wu et al. [36] classified existing works regarding the synergy of LLMs and evolutionary computation into two main branches: LLMs-based search operators [37, 38] and LLMs-based algorithm automation [39, 40, 41]. Our study can be roughly classified into the first category. In [42], LLMs are employed as crossover operators to derive new solutions from parental inputs. Brownlee et al. [43] also presented LLMs effectively functioning as mutation operators that enhance the search process. Liu et al. [44] introduced a novel framework known as LLMs-driven EA (LMEA), which utilizes LLMs for both crossover and mutation operations. Furthermore, LLMs-based search operators were adapted for multi-objective scenarios by segmenting traditional optimization tasks into sub-problems [45]. The related application can be found in [46], where protein properties are optimized with LLMs-enhanced evolutionary operators.
@@ II-C LLMs on Graph-Structured Problems
The recent advancements in LLMs have led to significant progress across diverse fields including sentiment analysis [47], optimization [48], and social sciences [49]. This raises the question of whether the capabilities of LLMs can be extended to graph-related tasks [50, 51]. However, as indicated by studies such as [52, 53], LLMs struggle with graph-structured data and often fail to achieve satisfactory results on fundamental tasks, especially as graph complexity increases.
Recent studies have explored the application of MLLMs to graph-related tasks, primarily as benchmarks to evaluate their visual reasoning capabilities. Wei et al. [54] combined visual and textual representations of graphs to enable MLLMs to perform structured reasoning, while VisionGraph [55] relied exclusively on visual inputs to assess how well these models interpret and reason over graphical structures. Beyond basic graph reasoning tasks, there are also some applications of MLLMs for combinatorial optimization. Huang et al. [56] proposed a multimodal approach that combines visual and textual inputs to solve the traveling salesman problem (TSP), whereas Elhenawy et al. [57] focused solely on visual representations to address routing tasks using MLLMs. Although both studies represent pioneering efforts in applying MLLMs to combinatorial problems, they do not incorporate evolutionary optimization frameworks and are limited to relatively small-scale networks with fewer than 200 nodes.
Our work is, to the best of our knowledge, the first to integrate MLLMs as evolutionary optimizers for solving large-scale combinatorial problems, with a particular focus on scalability and structural understanding of real-world complex networks. We aim to explore this untapped synergy by investigating how MLLMs can facilitate structure-aware evolutionary optimization in complex network scenarios, using influence maximization as a representative case study.
@@ III-B Graph Sparsification
Consider a graph $G=(V,E)$ with a set of communities $\widetilde{\mathcal{C}}=\{\mathcal{C}_{1},\mathcal{C}_{2},\ldots,\mathcal{C}_{m}\}$, where each $\mathcal{C}_{i}\subset V$ denotes a cluster of nodes. The simplified network should retain key nodes and connections, preserving the core community structure of the original network. The proposed sparsification process is as follows (see Algorithm 1 for the pseudocode): We begin by determining the number of selected nodes from each community, that is
@@
$V^{\prime}=\bigcup_{i=1}^{m}\{v_{j}\in\mathcal{C}_{i}\mid b(v_{j})\geq b(v_{l}),\forall v_{l}\in\mathcal{C}_{i}\setminus V^{\prime}\},$ (5)
where $b(v)$ refers to the betweenness value of node $v$ and $|V^{\prime}\cap\mathcal{C}_{i}|=|\mathcal{C}^{\prime}_{i}|$.
@@
In addition, isolated nodes will also be removed, and only the largest connected component is retained. The node and edge sets of refined network are updated as:
$V^{\prime}$ $=\{v\in V^{\prime}\mid d(v)>0\},$ (7)
$E^{\prime}$ $=\{(u,v)\in E^{\prime}\mid u,v\in V^{\prime}\}.$
@@
2:Identify the community structure $\widetilde{\mathcal{C}}=\{\mathcal{C}_{1},\mathcal{C}_{2},\ldots,\mathcal{C}_{m}\}$
3:for each community $\mathcal{C}_{i}$ in $\widetilde{\mathcal{C}}$ do
@@
8:Construct $V^{\prime}\leftarrow\bigcup_{i=1}^{m}V^{\prime}_{i}$
[…]
@@ III-C Community Merging
[…]
@@
    Size Identification: Identify the smallest community $\mathcal{C}_{\text{min}}$, where $|\mathcal{C}_{\text{min}}|=\min\{|\mathcal{C}_{i}|\}$ for all $i\in[1,m]$.
@@
    Connectivity Analysis: For each edge $(v_{i},v_{j})\in E$, increment an edge count between communities for edges where $v_{i}\in\mathcal{C}_{\text{min}}$ and $v_{j}\in\mathcal{C}_{k}\in\widetilde{\mathcal{C^{\prime}}}$, and vice versa. This edge count $e(\mathcal{C}_{\text{min}},\mathcal{C}_{k})$ measures the connectivity between $\mathcal{C}_{\text{min}}$ and the community $\mathcal{C}_{k}$.
@@
    Determination of Closest Community: Identify the community $\mathcal{C}_{\text{close}}$ with the maximum edge count to $\mathcal{C}_{\text{min}}$, i.e.,
    $\mathcal{C}_{\text{close}}=\arg\max_{\mathcal{C}_{k}\in\widetilde{\mathcal{C^{\prime}}}}e(\mathcal{C}_{\text{min}},\mathcal{C}_{k}).$
After each merging operation, the community structure $\widetilde{\mathcal{C^{\prime}}}$ is updated, and the process repeats until the number of communities equals $N_{\mathcal{C}}^{*}$. The process of community merging can be found in Algorithm 2.
@@
5:  for each community $\mathcal{C}_{k}\in\widetilde{\mathcal{C^{\prime}}}$ such that $\mathcal{C}_{k}\neq\mathcal{C}_{\min}$ do
@@
13:  Identify $\mathcal{C}_{\text{close}}$ to $\mathcal{C}_{\min}$ and merge them
@@ III-D MLLM-based Evolutionary Optimizer
When solving graph-structured combinatorial problems, traditional EAs typically utilize list-based encodings, consisting of sequences of node indices. While this approach is straightforward, it has some inherent drawbacks. Crossover and mutation operations could be compromised due to the absence of topological awareness concerning the network’s structure.
@@
[…]
@@
7:  $\text{Fitness}\leftarrow$ Evaluate the fitness of each node set in $\mathcal{P}$
8:  $\text{Parents}\leftarrow$ Select individuals from $\mathcal{P}$ based on fitness for reproduction
9:  $\text{Offspring}\leftarrow$ Generate new solutions by applying MLLM-based crossover to images of parent solutions
10:  Apply MLLM-based mutation to images of offspring solution to alter node sets
@@
[…]
@@
Figure 1: The diagram of MLLM-driven VEO and the comparison with traditional evolutionary optimization, illustrated with influence maximization. (a): The original network presents a graph with nodes and edges, where two solutions (seed sets) are shown: Solution 1 [A, C, P] and Solution 2 [A, E, H], representing selected nodes to maximize influence spread. (b): Traditional encoding and reproduction represent solutions as index lists. However, the crossover process often generates invalid solutions, such as duplicate nodes (e.g., [A, A, P]) or suboptimal combinations (e.g., [A, E, P] where nodes A and E are low-degree). (c): Our scheme encodes the solution in the form of an image where seed nodes of the solution are distinctly highlighted. Crossover intelligently combines nodes from parent solutions to produce valid offspring by filtering out low-degree nodes (e.g., nodes A and P), while mutation replaces ineffective nodes with more suitable ones to enhance influence spread (e.g., replacing node E with node F). This approach ensures validity and better solution quality. (d): Prompt for reproduction provides instructions for an MLLM: in crossover, the MLLM examines network images and suggests an optimal combination of seed nodes; in mutation, it identifies ineffective nodes for replacement by removing nodes with minimal influence and proposing better non-seed nodes.
TABLE I: Structure of prompts for MLLM-based evolutionary operators of different phases. The prompt begins with context-setting, which introduces the input information and clarifies the role of the agents. The latter part of the prompt specifies the desired output format and includes necessary guidance or restrictions.
@@
To overcome these challenges, we propose a new context-aware framework VEO that incorporates visual encoding and direct interaction with network images by utilizing the capabilities of MLLMs. This method transforms the solution into an image format, enhancing machines’ understanding of how specific nodes influence particular tasks. The diagram of our method is shown in Figure 1 and the process of MLLM-based evolutionary optimization is in Algorithm 3.
@@
which produces an image of the graph $G$ using the layout $\Theta$ via visualization software. In contrast, to encode specific candidate solutions, we define the function
@@
MLLM-based Mutation: A mutation operator is similarly applied on the network, yielding: